OpenStudy (anonymous):

Improper Integrals: Evaluate the integral from 0 to 1 of x^-1 dx, or show that it is divergent?

OpenStudy (anonymous):

\[=\int\limits_{0}^{1}dx/x=lnx\] | 0 to 1 ln(1) = 0, ln (0) -> -infinity => divergent

OpenStudy (anonymous):

the integrand is raised to the negative one, in other words int from 0 to 1 of 1/ sqrt x dx

OpenStudy (anonymous):

negative 1 is not not negative 1/2.... what is suppose to be...? \[1/\sqrt{x}\] or 1/x ...?

OpenStudy (anonymous):

Sorry for the confusion. The problem is int from 0 to 1 of dx/ sqrt x.

OpenStudy (anonymous):

OK... then: \[\int\limits_{0}^{1}dx/\sqrt{x}=2 * \sqrt{x}\] |0 to 1 =2 right?